OpenStudy (anonymous):

The length of a rectangle is 2 cm greater than the width. The area is 80 cm. Find the length and width.

OpenStudy (anonymous):

let the width by x cm length= (x+2)cm area = (x+2) times (x) = 80

OpenStudy (anonymous):

fLength 10cm Width 8 cm Area = Length x Width 80cm(cubed) = 10cm x 8cm

OpenStudy (anonymous):

\[(x+2) \times x = 80\] \[x ^{2} + 2x-80 = 0\]

OpenStudy (anonymous):

solve the equation and get x

OpenStudy (anonymous):

or you could do trial and error like fosterthepeople

OpenStudy (anonymous):

but that would be hard for a larger area in another similar question

OpenStudy (anonymous):

how did you get 80?

OpenStudy (anonymous):

80 is the area

OpenStudy (anonymous):

length x breadth = area

OpenStudy (anonymous):

did you see how i did it up there?

OpenStudy (anonymous):

Yup, I saw it.

OpenStudy (anonymous):

Irkiz and Fosterthepeople THANKS GUYS! BIG HELP!

OpenStudy (anonymous):

no problem :P